# 1st Wanqi Sun, 3rd Chenghua Duan School of Electronic, Electrical and Communication Engineering

### v1

@@ Abstract
The variational quantum eigensolver, as a hybrid quantum-classical algorithm, is widely applied in quantum chemistry simulations, especially in computing the electronic structure of complex molecular systems. However, on existing noisy intermediate-scale quantum devices, factors such as quantum decoherence, measurement errors, and gate operation imprecisions are unavoidable. To overcome these challenges, this study proposes an efficient noise-mitigating variational quantum eigensolver for accurate computation of molecular ground state energies in noisy environments. We design the quantum circuit with reference to the structure of matrix product states and utilize it to pre-train the circuit parameters, which ensures circuit stability and mitigates fluctuations caused by initialization. We also employ zero-noise extrapolation to mitigate quantum noise and combine it with neural networks to improve the accuracy of the noise-fitting function, which significantly eliminates noise interference. Furthermore, we implement intelligent grouping measurements of Hamiltonian Pauli strings, which not only reduces measurement errors but also improves sampling efficiency. We conduct experiments to solve the ground state energy of the $H_{4}$ molecule on the Mindquantum quantum computing platform, and the results demonstrate that our algorithm can constrain noise errors within the range of $\mathcal{O}(10^{-2})\sim\mathcal{O}(10^{-1})$, outperforming mainstream variational quantum eigensolver methods. This work provides a new strategy for high-precision quantum chemistry calculations on near-term noisy quantum hardware.
@@ I Introduction
In this context, the Variational Quantum Eigensolver (VQE), as a hybrid quantum-classical algorithm, has become an important tool for achieving this goal [3, 4]. Matrix product states (MPS) are also widely applied as the VQE because their one-dimensional chain structure is effective in capturing the local features of quantum states [5, 6, 7]. The primary objective of the VQE is to determine the ground state energy and its corresponding quantum state for the Hamiltonian $H$ of a closed quantum system, which is crucial for predicting various chemical properties including reaction rates and stable molecular configurations [8, 9]. In 2014, Peruzzo et al. first applied the Variational Quantum Eigensolver (VQE) combined with unitary coupled cluster theory to quantum chemistry simulations [10, 11], successfully determining the ground state energy of the He-H+ system.
@@
To address the noise challenges in current noisy quantum devices, we propose an innovative variational quantum eigensolver solution. The main contributions of our work are as follows:
@@
    Design quantum circuit structures based on MPS, leveraging the chain structure of MPS to capture local entanglement while ensuring that the circuit depth remains sufficiently shallow.
@@
    Implement grouped measurements of Hamiltonian Pauli strings, which reduces the number of sampling and mitigates measurement noise.
Experiments are conducted using the quantum computing framework MindQuantum, where the noise environment includes depolarizing noise, thermal relaxation noise, and bit-flip noise. The experimental results demonstrate that our algorithm can limit noise errors within $\mathcal{O}(10^{-2})\sim\mathcal{O}(10^{-1})$, surpassing current mainstream variational eigensolver algorithms.
Figure 1: Schematic Representation of MPS and the Algorithm. (a) depicts a 5-qubit quantum state |$\ket{\psi}$ in the MPS form, where the $n$th local tensor $A^{[n]}$ corresponds to the $n$th qubit. The MPS shown below has been gauged into a center-orthogonal form, with $n_{c}$ as the orthogonal center. (b) describes the eigenvalue computation process when the MPS is employed as the trial state. (c) illustrates how the center-orthogonal MPS can reduce the computational cost of eigenvalue calculations, requiring only the contraction of tensors from the $i$th to the $n_{c}$th site with the Hamiltonian coefficient tensor. (d) outlines our MPS-VQE algorithm. The pre-trained MPS parameters serve as the initial values for the quantum circuit parameters $\theta$, which are then further optimized. The circuit folding is employed to obtain expectation values with different noise levels, and then fits the values to extrapolate the noise-free expectation value.
@@ II-A Variational Quantum Eigensolver Framework
The VQE is a hybrid quantum-classical algorithm designed to compute the ground state energy of quantum systems. The main implementation of VQE involves preparing a parameterized trial wavefunction $|\psi(\boldsymbol{\theta})\rangle$ on a quantum device, then combining it with classical machine learning optimization algorithms such as gradient descent to continuously adjust the parameters $\boldsymbol{\theta}$ until the expectation value $\langle\psi(\boldsymbol{\theta})|H|\psi(\boldsymbol{\theta})\rangle$ is minimized. Mathematically as
@@
where $c_{i}$ is real coefficient and $H_{i}$ denotes the tensor product of Pauli operators, such as $H_{i}=\sigma_{i_{1}}\otimes\sigma_{i_{2}}\otimes\cdots\otimes\sigma_{i_{n}}$, $\sigma_{i_{k}}\in\{I,X,Y,Z\}.$
@@
    Wavefunction Preprocessing: Gauge the wavefunction |$\ket{\psi}$ to the center-orthogonal form to enhance computational stability and reduce the complexity of the expectation computation
@@
    Pre-training: Pre-train the matrix product syate on classical computer as the quantum circuit initialization.
@@
    Quantum State Measurement: Combine neural networks with zero-noise extrapolation techniques to mitigate measurement errors. Perform grouped measurements of the Hamiltonian Pauli strings to reduce the number of measurements required.
@@ II-B Wavefunction Construction
$\varphi_{\scriptscriptstyle s_{0}s_{1}\cdots s_{N-1}}=\sum_{\scriptscriptstyle\alpha_{0},\cdots,\alpha_{N-2}}A^{\left[0\right]}_{\scriptscriptstyle s_{0},\alpha_{0}}A^{\left[1\right]}_{\scriptscriptstyle s_{1},\alpha_{0},\alpha_{1}}\cdots A^{\left[N-1\right]}_{\scriptscriptstyle s_{N-1},\alpha_{N-2}}.$ (4)
The $\{\left|s_{n}\right\rangle\}$ denotes a set of orthonormal computational bases. As shown in Figure 1 (a), $A^{\left[n\right]}_{s_{n},\alpha_{n-1},\alpha_{n}}$ is the local tensor corresponding to the $n$-th qubit, where $\{s_{n}\}$ denotes the physical indices of dimension $d$, and $\{\alpha_{n}\}$ are the virtual indices of dimension $\chi$.
Now, the eigenvalue problem has been transformed from Eq. (1) to $E_{0}=\min_{\boldsymbol{\theta}}\sum_{i}c_{i}\langle\psi|H_{i,i+1}|\psi\rangle$, where the parameters $\boldsymbol{\theta}$ equel to $\{A^{[n]}\}$.
@@ II-C Wavefunction Preprocessing
Another key advantage is that it can significantly simplify the computation of the energy. As shown in Fig. 1(b), if the local coupling is located to the left of the orthogonal center, i.e., $i<n_{c}$, according to the left-right orthogonality condition, the contraction of the tensor to the left of the $i$th tensor is equal to the identity, and the contraction of the tensor to the right of the orthogonal center $n_{c}$ is also equal to the identity. It is only required to contract the $i$th to $n_{c}$th tensor with the Hamiltonian coefficient tensor. The simplification of the computation is exactly similar when $i+1\geq n_{c}$.
@@ II-D Wavefunction Pre-training
It is found that even for the same quantum circuit structure, if the initial parameter values are different, the measurement results of the trained quantum circuit will be different. Therefore, we adopted the MPS pre-training to initialize the quantum circuit parameters, which can preserve the local features of the quantum state in advance and improve the accuracy of the final calculation. The objective function for MPS pre-training is Eq. (1) and the update rule for the $n$th local tensor is
$A^{[n]}\leftarrow A^{[n]}-\eta\frac{\partial\sum_{i}\bra{\psi}H_{i,i+1}\ket{\psi}}{\partial A^{[n]}},$ (5)
@@ II-E Designing Quantum Circuits for the Wavefunction
[…]
@@
$\mathcal{U}^{[n]\dagger}=(U(\theta_{n,l_{0}})\otimes U(\theta_{n,l_{1}}))\ CNOT\ (U(\theta_{n,r_{0}})\otimes U(\theta_{n,r_{1}})).$ At $n=0$, the single qubit unitary $\mathcal{U}^{[n]}$ is factorized as:
$\mathcal{U}^{[0]\dagger}=RZ(\theta_{0})RY(\theta_{1})RZ(\theta_{2}).$
This decomposition ensures that each qubit ends in a Z-axis rotation at the end of the circuit. The Pauli Z-gate with eigenvalues of 0 and 1, the eigenvectors are consistent with the computational basis. Therefore, the energy value can be quickly calculated directly by sampling the measured quantum state, which improves the computational efficiency.
@@ II-F Noise Error Mitigation
When executing algorithms on quantum computing devices, noise interference is inevitable. For gate operation noise and relaxation noise, we apply the zero-noise extrapolation technique to mitigate the errors at the algorithmic level. This process works in two steps:
@@
TABLE I: Comparison of circuit metrics for different VQE models
@@ III Experiments and Results
Table II summarizes our MPS-VQE circuit implementation as well as other models, including key metrics such as the number of qubits, total gates, barriers, and parameterized gates. The number of layers in our MPS-VQE circuit is set to 1.
@@
Results. The Full Configuration Interaction (FCI) benchmark value for this assumed $H_{4}$ model is -2.1664 Hartree. Table II presents the results of different VQE models in solving the ground state energy of the $H_{4}$ molecule under both ideal noiseless and noisy environments. The listed values are the best results obtained from 30 independent experiments. The comparison results of pre-training MPS to initialize the quantum circuit are shown in Figure 2.
Our proposed MPS-VQE circuit, with a relatively small number of gates, achieves calculation accuracy comparable to the Unitary Coupled Cluster Singles and Doubles (UCCSD) method, which has around 30 times more gates, in the ideal noiseless scenario. More importantly, our MPS-VQE demonstrates significantly enhanced noise tolerance compared to other VQE algorithms. This indicates that the MPS-VQE can more effectively mitigate the impact of noise interference in actual quantum hardware, thereby exhibiting greater robustness for practical applications.
@@ IV Conclusion and Discuss
This work proposes an innovative variational quantum eigensolver algorithm that can effectively mitigate the unavoidable noise in NISQ devices. We have introduced innovations in three aspects: circuit initialization, circuit structure design, and error mitigation strategies for both computational and measurement errors. The research findings demonstrate that our proposed MPS-VQE scheme can achieve the precision of large-scale complex circuits using a much more compact circuit, while exhibiting superior noise resilience compared to other VQE algorithms.
This study provides new approaches and methodologies for realizing high-precision quantum computing on medium-noise quantum hardware, which is of great significance in advancing the practical applications of quantum computing. Future work may further explore the application of this scheme to more complex quantum chemistry and materials science problems, as well as its broader adoption in a wider range of quantum algorithms, aiming to enhance the practical utility of quantum computation.

### v2

# Noise-Mitigated Variational Quantum Eigensolver with Pre-training and Zero-Noise Extrapolation

@@ Abstract
As a hybrid quantum-classical algorithm, the variational quantum eigensolver is widely applied in quantum chemistry simulations, especially in computing the electronic structure of complex molecular systems. However, on existing noisy intermediate-scale quantum devices, some factors such as quantum decoherence, measurement errors, and gate operation imprecisions are unavoidable. To overcome these challenges, this study proposes an efficient noise-mitigating variational quantum eigensolver for accurate computation of molecular ground state energies in noisy environments. We design the quantum circuit with reference to the structure of matrix product states and utilize it to pre-train the circuit parameters, which ensures circuit stability and mitigates fluctuations caused by initialization. We also employ zero-noise extrapolation to mitigate quantum noise and combine it with neural networks to improve the accuracy of the noise-fitting function, which significantly eliminates noise interference. Furthermore, we implement an intelligent grouping strategy for measuring Hamiltonian Pauli strings, which not only reduces measurement errors but also improves sampling efficiency. We perform numerical simulations to solve the ground state energy of the $H_{4}$ molecule by using MindSpore Quantum framework, and the results demonstrate that our algorithm can constrain noise errors within the range of $\mathcal{O}(10^{-2})\sim\mathcal{O}(10^{-1})$, outperforming mainstream variational quantum eigensolvers. This work provides a new strategy for high-precision quantum chemistry calculations on near-term noisy quantum hardware.
@@ I Introduction
In this context, the Variational Quantum Eigensolver (VQE), as a hybrid quantum-classical algorithm, has become an important tool for achieving this goal [3, 4]. Matrix product states (MPS) are also widely applied as the VQE because their one-dimensional chain structure is effective in capturing the local features of quantum states [5, 6, 7]. The primary objective of the VQE is to determine the ground state energy and corresponding quantum state of the Hamiltonian $H$ for a closed quantum system, which is crucial for predicting various chemical properties including reaction rates and stable molecular configurations [8, 9, 10]. In 2014, Peruzzo et al. first applied the Variational Quantum Eigensolver (VQE) combined with unitary coupled cluster theory to quantum chemistry simulations [11], successfully determining the ground state energy of the He-H+ system.
@@
To tackle the noise challenges in current quantum devices, we propose an innovative variational quantum eigensolver. The main contributions of our work are as follows:
@@
    Design one adjustable quantum circuit structure based on MPS, fully utilizing the chain structure of MPS to capture local entanglement, while considering a brick-wall structure with similar error scaling to ensure the circuit depth remains shallow.
@@
    Implement grouping measurements of Hamiltonian Pauli strings, which reduces the number of sampling and mitigates measurement noise.
The numerical simulations are mainly performed by using MindSpore Quantum [21] 0.9.11 framework, where the noise environment includes depolarizing noise, thermal relaxation noise, and bit-flip noise. The experimental results demonstrate that our algorithm can limit noise errors within $\mathcal{O}(10^{-2})\sim\mathcal{O}(10^{-1})$, surpassing mainstream variational eigensolvers.
Figure 1: Schematic Representation of MPS and the Algorithm. (a) depicts a 5-qubit quantum state |$\ket{\psi}$ in the MPS form, where the $n$th local tensor $A^{[n]}$ corresponds to the $n$th qubit. The MPS shown below has been gauged into a center-orthogonal form with $n_{c}$ as the orthogonal center. (b) describes the eigenvalue computation process when the MPS is employed as the trial state. (c) illustrates how the center-orthogonal MPS can reduce the computational cost of eigenvalue calculations, requiring only the contraction of tensors from the $i$th to the $n_{c}$th site with the Hamiltonian coefficient tensor. (d) outlines our MPS-VQE algorithm. The pre-trained MPS parameters serve as the initial values for the quantum circuit parameters $\theta$, which are then further optimized. The circuit folding is employed to obtain expectation values with different noise levels, and then fits the values to extrapolate the noise-free expectation value.
@@ II-A Variational Quantum Eigensolver Framework
The VQE is a hybrid quantum-classical algorithm designed to compute the ground state energy of quantum systems. The main implementation of VQE involves preparing a parameterized trial wavefunction $|\psi(\boldsymbol{\theta})\rangle$ on a quantum device, then combining it with classical machine learning optimization algorithms such as gradient descent to continuously adjust the parameters $\boldsymbol{\theta}$ until the expectation value $\langle\psi(\boldsymbol{\theta})|H|\psi(\boldsymbol{\theta})\rangle$ is minimized. Mathematically, this is expressed as:
@@
where $c_{i}$ is real coefficient and $H_{i}$ is the tensor product of Pauli operators, such as $H_{i}=\sigma_{i_{1}}\otimes\sigma_{i_{2}}\otimes\cdots\otimes\sigma_{i_{n}}$, $\sigma_{i_{k}}\in\{I,X,Y,Z\}.$
@@
    Wavefunction Preprocessing: Gauge the wavefunction |$\ket{\psi}$ to the center-orthogonal form to enhance computational stability and reduce the complexity of the expectation computation.
@@
    Pre-training: Pre-train the matrix product state on classical computer as the quantum circuit initialization.
@@
    Quantum State Measurement: Combine neural networks with zero-noise extrapolation techniques to mitigate measurement errors. Perform grouping measurements of the Hamiltonian Pauli strings to reduce the number of measurements required.
@@ II-B Wavefunction Construction
$\small\varphi_{\scriptscriptstyle s_{0}s_{1}\cdots s_{N-1}}=\sum_{\scriptscriptstyle\alpha_{0},\cdots,\alpha_{N-2}}A^{\left[0\right]}_{\scriptscriptstyle s_{0},\alpha_{0}}A^{\left[1\right]}_{\scriptscriptstyle s_{1},\alpha_{0},\alpha_{1}}\cdots A^{\left[N-1\right]}_{\scriptscriptstyle s_{N-1},\alpha_{N-2}}.$ (4)
The $\{\left|s_{n}\right\rangle\}$ denotes a set of orthonormal computational bases. As shown in Figure 1 (a), $A^{\left[n\right]}_{s_{n},\alpha_{n-1},\alpha_{n}}$ is the local tensor corresponding to the $n$-th qubit, where $\{s_{n}\}$ is a set of the physical indices of dimension $d$, and $\{\alpha_{n}\}$ is a set of the virtual indices of dimension $\chi$.
Now, the eigenvalue problem has been transformed from Eq. (1) to $E_{0}=\min_{\boldsymbol{\theta}}\sum_{i}c_{i}\langle\psi|H_{i,i+1}|\psi\rangle$, where the parameter $\boldsymbol{\theta}$ refers to $\{A^{[n]}\}$.
@@ II-C Wavefunction Preprocessing
Another key advantage is that it can significantly simplify the computation of the energy. As shown in Fig. 1(b), if the local coupling is located to the left of the orthogonal center, i.e., $i<n_{c}$, according to the left-right orthogonality condition, the contraction of the tensor to the left of the $i$th tensor is equal to the identity, and the contraction of the tensor to the right of the orthogonal center $n_{c}$ is also equal to the identity. It is only required to contract the $i$th tensor to $n_{c}$th one with the Hamiltonian coefficient tensor. The simplification of the computation is exactly similar when $i+1\geq n_{c}$.
@@ II-D Wavefunction Pre-training
It is found that even for the same quantum circuit structure, if the initial parameter values are different, the measurement results of the trained quantum circuit will be different. Therefore, we adopted the MPS pre-training to initialize the quantum circuit parameters, which can preserve the local features of the quantum state in advance and improve the accuracy of the final calculation. The objective function for MPS pre-training is defined in Eq. (1) and the update rule for the $n$th local tensor is defined as follows.
$A^{[n]}\leftarrow A^{[n]}-\eta\frac{\partial\sum_{i}\bra{\psi}H_{i,i+1}\ket{\psi}}{\partial A^{[n]}}$ (5)
@@ II-E Designing Quantum Circuits for the Wavefunction
In this paper, with reference to the structure of MPS for the circuit framework of wave functions, the $n$th local tensor $A^{[n]}$ simulates the quantum state evolution and entanglement on the $n$th qubit. The quantum circuit framework is shown in Figure 1 (c), where the local tensor $A^{[n]}$ is mapped to a parameterized quantum circuit $\mathcal{U}^{[n]\dagger}$. When constructing the wave function $|\psi\rangle$, the operators $\{\mathcal{U}^{[n]\dagger}\}$ act sequentially on the initial Hartree-Fock state $|\psi_{\mathrm{HF}}\rangle$:
@@
$\small\mathcal{U}^{[n]\dagger}=(U(\theta_{n,l_{0}})\otimes U(\theta_{n,l_{1}}))\ CNOT\ (U(\theta_{n,r_{0}})\otimes U(\theta_{n,r_{1}})),$ (7)
where $U(\theta_{*})$ is a single qubit unitary that can be further decomposed as:
$U(\theta_{*})=RZ(\theta_{0})RY(\theta_{1})RZ(\theta_{2}).$ (8)
For $n=0$, $\mathcal{U}^{[0]}$ is a single qubit unitary, which is expressed as Eq. (8).
This decomposition ensures that each qubit ends in a Z-axis rotation at the end of the circuit. The Pauli Z-gate has eigenvalues of 0 and 1, and its eigenvectors are consistent with the computational basis. Therefore, the energy value can be quickly calculated directly by sampling the measured quantum state, which improves the computational efficiency.
@@ II-F Noise Error Mitigation
When executing algorithms on quantum computing devices, noise interference is inevitable. For gate operation noise and relaxation noise, we apply the zero-noise extrapolation technique to mitigate the errors at the algorithmic level. This process includes two steps:
@@
TABLE I: Comparison of circuit metrics for different VQE models
@@ III Experimental Results
Table II summarizes our MPS-VQE circuit implementation and other models, including key metrics such as the number of qubits, total gates, barriers, and parameterized gates. The number of layers in our MPS-VQE circuit is set to 1.
@@
Experimental Results. The Full Configuration Interaction (FCI) benchmark value for this assumed $H_{4}$ model is -2.1664 Hartree. Table II presents the results of different VQE models in solving the ground state energy of the $H_{4}$ molecule under both ideal noiseless and noisy environments. The listed values are the best results obtained from 30 independent experiments. The comparison results of pre-training MPS to initialize the quantum circuit are shown in Figure 2.
Under noiseless conditions, our proposed MPS-VQE circuit, with fewer gates, achieves similar calculation accuracy to the Unitary Coupled Cluster Singles and Doubles (UCCSD) method, which uses approximately 30 times the number of gates. More importantly, our MPS-VQE demonstrates significantly enhanced noise tolerance compared to other VQE algorithms. This indicates that the MPS-VQE can more effectively mitigate the impact of noise interference in actual quantum hardware, thereby exhibiting greater robustness for practical applications.
@@ IV Conclusion and Discuss
This work proposes an innovative variational quantum eigensolver that can effectively mitigate the unavoidable noise in NISQ devices. We have introduced innovations in three aspects: circuit initialization, circuit structure design, and error mitigation strategies for both computational and measurement errors. The research findings demonstrate that our proposed MPS-VQE scheme can achieve the precision of large-scale complex circuits using a much more compact circuit, while exhibiting superior noise resilience compared to other VQEs.
This study provides one new approach for realizing high-precision quantum computing on medium-noise quantum hardware, which is of great significance in advancing the practical applications of quantum computing. Future work may further explore the application of this scheme to more complex quantum chemistry and materials science problems, as well as its generalization to more quantum algorithms, aiming to enhance the practical utility of quantum computation.
Acknowledgment
Thanks for the support provided by MindSpore Community.